OpenStudy (anonymous):

What is the value of the expression 3 – a – b when a = 15 and b = –3? A. –21 B. –15 C. –9 D. 21

OpenStudy (anonymous):

3-15-(-3)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

it would be B. -15 because 3-15= -12 -(-3)(You would add them together) would give you -15

OpenStudy (anonymous):

You would rewrite the expression like this 3-15-(-3), then do 3-15= -12, and then you do -12-(-3)= -9.